OpenStudy (anonymous):

Evualte the integral integration{(1+sqrt3)^3}/sqrtx evualuate with proper substitution answer this what should i have to do with this?

OpenStudy (zarkon):

\[\int\frac{(1+\sqrt{3})^3}{\sqrt{x}}dx\] this?

OpenStudy (anonymous):

the top is just a constant so u can take it outside the integral and the bottom is x^(-1/2)

OpenStudy (anonymous):

Or to be clear, the integral is for x^(-1/2)

OpenStudy (anonymous):

but which will u ?

OpenStudy (anonymous):

U don't need substitution for that.

OpenStudy (zarkon):

if your problem is as I typed it above there is no need for substitution.